Toni said "i bet you cant guess my numbers! the first is five more than the second. if you add the first to three times the second you get 29." what are toni's numbers?
Answer by mananth(16946)

let the second number be x
first number = x+5
..
3x+x+5=29
4x+5=29
4x=24
x=6 first number
second number = 6+5=11